# Super-intelligence or Superstition? Exploring Psychological Factors Underlying Unwarranted Belief in AI Predictions

### v2

@@ Abstract
This study investigates psychological factors influencing belief in AI predictions about personal behavior, comparing it to belief in astrology and personality-based predictions. Through an experiment with 238 participants, we examined how cognitive style, paranormal beliefs, AI attitudes, personality traits, and other factors affect perceived validity, reliability, usefulness, and personalization of predictions from different sources. Our findings reveal that belief in AI predictions is positively correlated with belief in predictions based on astrology and personality psychology. Notably, paranormal beliefs and positive AI attitudes significantly increased perceived validity, reliability, usefulness, and personalization of AI predictions. Conscientiousness was negatively correlated with belief in predictions across all sources, and interest in the prediction topic increased believability across predictions. Surprisingly, cognitive style did not significantly influence belief in predictions. These results highlight the "rational superstition" phenomenon in AI, where belief is driven more by mental heuristics and intuition than critical evaluation. We discuss implications for designing AI systems and communication strategies that foster appropriate trust and skepticism. This research contributes to our understanding of the psychology of human-AI interaction and offers insights for the design and deployment of AI systems.
Keywords AI  $\cdot$ Large Language Model  $\cdot$ Cognitive Biases  $\cdot$ Mental Models  $\cdot$ Prediction  $\cdot$ Belief
@@ 1 Introduction
Technology is often associated with scientific advancement and frequently viewed in opposition to superstition. However, the recent rapid advancements in artificial intelligence (AI) have given rise to quasi-religious perspectives among some leaders in the tech industry [1]. In fact, as generative AI and large language models (LLMs) continue to evolve, a contemporary critique posits that the narrative surrounding their development bears striking resemblances to religious discourse [2]. Promoting concepts such as superintelligence and sentience, prominent figures in the tech industry have been observed making prophetic claims regarding AI’s potential to either salvage or obliterate humanity [3, 4]. These claims are often based on personal speculation and subjective views rather than rigorous scientific research. Nonetheless, empirical research has demonstrated that these popular narratives are mirrored in public perceptions of AI, frequently manifesting as exaggerated utopian or dystopian visions [5, 6]. This intersection of technological advancement and quasi-religious perspective presents a complex landscape for scholars to navigate, as it blurs the boundaries between scientific progress and speculation.
While AI has been developed through predictive algorithms and statistical pattern recognition, exemplifying advancements in mathematical and rational thinking, public perception of AI may draw unexpected parallels with astrology [7, 8, 9]. Despite astrology representing an opposing epistemology based on pseudoscientific methodologies antithetical to empirical reasoning, both AI and astrology function as prediction-generating mechanisms helping people make sense of the uncertainty in the world [7, 8]. Anthropologist Christophe Lazaro suggests that AI, powered by data, has become a new mode of speculative thinking, with algorithms taking on the role of oracles in contemporary societies, framing AI as "artificial divination" [8, p. 128].
On the other hand, AI, unlike astrology, presents an aura of scientific legitimacy, backed by a vast body of scientific research in computer science, mathematics, cognitive science, and more. This foundation contributes to the belief that AI systems are rational and objective. However, studies have shown that algorithms could produce unpredictable outcomes, biased results, and inherit implicit societal biases [10, 11, 12]. Moreover, despite ongoing efforts to increase transparency, much of the inner workings of AI systems remain "black boxes" [13, 14, 15, 16]. Such misconceptions regarding AI may engender a form of naive technological optimism and irrational trust in the system, which Wilson refers to as "rational superstition" [17]. Scholars have investigated the concept of algorithmic appreciation—the propensity to rely on algorithms for decision-making processes—which may lead to an overreliance on AI in critical decision-making contexts [18, 19].
Popular narratives that over-inflate the promises of AI and its capabilities is particularly alarming when considering the impact it has on mental models and perception of AI systems [6]. According to the dual process theory in psychology, people engage in two types of thinking—fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Incomplete or incorrect mental models shaped by speculative claims about AI rather than scientific evidence can become a mental heuristic that triggers people’s System 1 thinking rather than engaging their System 2 thinking, blurring the lines between speculation and informed reasoning [22]. This, in turn, can lead to miscalibration of trust and over-reliance on AI, which drives ineffective use of AI [18, 19, 23].
For instance, consider an AI system that claims to give you personalized product recommendations based on your social media usage and purchase history. A user with an incorrect mental model of the system’s capabilities and limitations may hastily rely on its recommendations without considering other options. On the flip side, a user who has developed an overly skeptical mental model of AI may reject helpful recommendations and miss critical opportunities. Therefore, without ongoing effort to establish appropriate mental models, users can become more prone to manipulation and potential negative outcomes.
While belief in AI supremacy is growing despite its demonstrated limitations and biases, our study presents the first empirical investigation into the phenomenon of "rational superstition" [17] in AI. To do so, we examine how people’s perception and trust in AI predictions may correlate with other irrational behaviors, such as belief in astrology. This research question is particularly timely and relevant given the recent surge in popularity of astrology, especially among Millennials and Gen Z [24, 25], despite the lack of empirical evidence supporting its validity [26, 27]. A recent consumer survey of over 2000 people in the US revealed that about a third of Millennials and a quarter of Gen Z have made financial decisions based on their horoscope [28].
To address this research question, we conducted an experiment involving 238 participants who were presented with fictitious predictions from three sources (AI, astrology, and personality psychology) and rated their perceived validity, reliability, usefulness, and personalization. To create a reasonable basis for the predictions, participants answered questionnaires about astrology and personality and engaged in a simulated investment game that appeared to analyze their interactions and predict their future investment behavior. Participants were randomly assigned to either positive (N = 119) or negative (N = 119) prediction groups, receiving forecasts about their future investment behavior and outcomes accordingly.
@@
Our study also explores how cognitive style, paranormal beliefs, gullibility, trust in AI, and personality traits may influence belief in predictions based on AI, astrology, and personality psychology. This investigation builds upon previous work examining how user characteristics such as personality traits influence the perception of AI-generated advice [29] and trust in AI [30]. Additionally, we consider the relationship between cognitive style and paranormal beliefs, including astrology [31, 32, 33, 34, 35], as well as how cognitive reflection affects social media usage and the sharing of (mis)information and conspiracy theories [36, 37].
Furthermore, we explore how demographic factors such as age, gender, and education level, as well as prior experience in AI, astrology, and personality psychology, and the level of interest in the topic influence belief in the predictions. By examining these various factors, our research aims to provide a comprehensive understanding of the "rational superstition" phenomenon in AI and its implications for decision-making in an increasingly AI-driven world.
Based on our research questions, we tested the following hypotheses:
@@
  • H2: Individuals with higher cognitive reflection and need for cognition will find AI predictions more credible and reliable than astrology- or personality-based predictions.
  • H3: Demographic factors and personal attributes (cognitive reflection, trust in AI, paranormal beliefs, gullibility, personality, level of interest, and prior experience) will moderate the perceived validity, reliability, usefulness, and personalization of predictions from different sources.
The study’s findings contribute valuable insights to the ongoing discourse on AI ethics, transparency, and responsible implementation. By shedding light on the factors that influence trust and belief in AI predictions, we aim to inform policymakers, AI developers, and educators about potential pitfalls and areas for improvement in AI literacy and public engagement with AI technologies.
@@ 2 Results
To test the first hypothesis (H1), a multiple linear regression model was applied to the data in wide format, with believability score for AI predictions as the outcome variable and believability scores for astrology- and personality-based predictions as predictor variables, along with control variables (cognitive style, paranormal beliefs, gullibility, AI attitude/trust in AI, big five personality, familiarity with the prediction sources (AI, astrology, personality), interest in topic of prediction, age, gender, education level). The scores across the four subscales (perceived validity, reliability, usefulness, and personalization) were averaged as overall "believability" scores to provide a straightforward test of the hypothesis.
The multiple linear regression analysis explained a significant proportion of the variance in the AI overall score ($R^{2}=0.7606$, Adjusted $R^{2}=0.7337$, $F(24,213)=28.2$, $p<0.001$). Significant predictors included the zodiac overall score (Estimate = 0.3119, $p<0.001$) and the personality overall score (Estimate = 0.4585, $p<0.001$), supporting the hypothesis that belief in astrology and personality-based predictions is positively associated with belief in AI predictions. This association can be visually observed in the scatterplots in Figure 1.
Figure 1: Relationship between belief in AI predictions and astrology predictions (left) and belief in AI predictions and personality predictions (right). The scatterplots show individual scores, with a linear regression line indicating a positive correlation. The shaded area represents the 95% confidence interval for the regression line.
@@
To examine the influence of moderating factors including cognitive style (H2), paranormal beliefs, trust in AI, personality traits, demographic factors, etc. (H3), a mixed-effects model was fitted to the data in a nested long format, with prediction source (AI, astrology, personality) and believability subscale (perceived validity, reliability, usefulness, and personalization) as categorical predictors, and subscale score as the outcome variable. This approach allowed examination of the main effects and moderating effects of psychological and cognitive factors on belief in AI predictions.
@@
Based on the mixed effects model analysis, the baseline subscale score, associated with "AI" prophecy source, "Validity" subscale, "Positive" prophecy group, "Female" gender, and "Bachelor" education level, was 5.10 ($p<0.001$) on a 7-point Likert scale. The main effects of prophecy source on believability of AI predictions showed that on average, perceived validity of astrology-based predictions was 0.66 points lower compared to AI predictions (95% CI [-0.93, -0.39], p < 0.001). The difference was not statistically significant for personality-based predictions (0.07, 95% CI [-0.18, 0.32], p = 0.593). These results suggest that people generally found fictitious predictions from AI, astrology, and personality convincing.
@@
There were significant main effects of subscales on the perception of AI predictions. On average, perceived personalization was rated 0.23 point higher than perceived validity (95% CI [0.08, 0.38], p = 0.002). On the other hand, perceived reliability was rated 0.91 points lower than perceived validity (95% CI [-1.13, -0.68], p < 0.001), and perceived usefulness was rated 0.63 points lower than perceived validity (95% CI [-0.83, -0.44], p < 0.001). This relationship between the subscales by prediction source can be visualized in Figure 2. The interaction between subscales and prophecy source showed that the main effects of subscales were largely consistent across AI, astrology, and personality, with some subscales being more pronounced for astrology. Perceived personalization was further increased by 0.23 points (95% CI [0.02, 0.44], p = 0.035) and perceived reliability further decreased by 0.31 (95% CI [-0.62, -0.00], p = 0.048) for astrology-based predictions, compared to the AI baseline. Other interaction terms were not significant, suggesting that the effects are not significantly different from the main effects.
@@
Furthermore, we found that compared to the baseline “Positive” prediction group, the “Negative” prediction group was associated with a decrease in perceived validity by 1.19 points (95% CI [-1.53, -0.85], p < 0.001). We found no significant interactions between prophecy group and prophecy source, suggesting that this effect was consistent across AI, astrology, and personality. This suggests that while people generally found fictitious predictions believable, it was not the case for those who received negative predictions.
@@
Figure 2: Boxplot showing the distribution of subscale scores by prophecy source and subscale. The boxes represent the interquartile range, with the means indicated in green triangles and notches representing the 95% confidence intervals for the medians.
@@
2.3 There is no evidence of correlation between belief in predictions and cognitive style.
@@
Based on our hypothesis (H2), we expected to see either a significant positive association between cognitive style and belief in AI predictions, significant negative interactions between cognitive style and other prediction sources (astrology, personality), or both. However, our results showed that the composite cognitive score, as measured as the composite of the performance-based Cognitive Reflection Test (CRT-2) [38] and the preference-based Need for Cognition (NFC-6) [39], did not significantly increase the perceived validity of AI predictions. The estimated increase in perceived validity was 0.13 points with one point increase in the composite cognitive score, but this effect was not statistically significant (95% CI [-0.01, 0.26], p = 0.065). The composite cognitive score ranged from -4.29 to 2.61 (Mean = 0.0, SD = 1.44).
However, we found some significant negative interactions between cognitive score and the subscales. Compared to perceived validity, higher composite cognitive score was associated with a decrease in perceived reliability (-0.11, 95% CI [-0.20, -0.02], p = 0.021) and perceived usefulness (-0.12, 95% CI [-0.20, -0.04], p = 0.004), making the main effect less positive. The interaction with perceived personalization was not statistically significant (-0.05, 95% CI [-0.11, 0.01], p = 0.092).
The interactions between cognitive style and prediction sources showed that while there was some negative interaction for astrology (-0.10, 95% CI [-0.21, 0.01], p = 0.079) and personality (-0.07, 95% CI [-0.18, 0.03], p = 0.170), the effects were not statistically significant. This suggests a lack of evidence that cognitive style is an influential factor in how people perceive predictions based on AI, astrology, and personality. Three-way interactions between subscales, prophecy source, and composite cognitive score did not lead to significant results.
@@
2.4 Higher paranormal beliefs increases perceived validity, reliability, usefulness, and personalization of AI predictions.
@@
One of the most interesting findings was that having paranormal beliefs was positively associated with belief in AI predictions. The results from the mixed effects model showed that the paranormal beliefs score, as measured by a shortened version of the Revised Paranormal Belief Scale (R-PBS) [40], significantly increased perceived validity of AI predictions. With each point increase in the paranormal beliefs scale, perceived validity of AI predictions increased by an average of 0.02 points (95% CI [0.01, 0.03], p = 0.001) on a 7-point scale. The paranormal score, which was centered for the analysis, originally ranged from 15 to 95 (Mean = 45.6, SD = 20.08).
@@
The interaction between paranormal beliefs and subscales showed that the effect was stronger for perceived reliability (0.01, 95% CI [0.00, 0.01], p = 0.020) and perceived usefulness (0.01, 95% CI [0.01, 0.02], p < 0.001), both being statistically significant. There was no significant interaction effect for perceived personalization (0.00, 95% CI [-0.00, 0.01], p = 0.427), suggesting that the effect did not differ significantly from perceived validity.
@@
Paranormal beliefs score was an even stronger predictor for belief in astrology-based predictions, perhaps less surprisingly given that the scale includes questions around astrology (see Section 7.3). Each standard deviation increase in paranormal beliefs led to an additional 0.01 point increase in perceived validity (95% CI [0.01, 0.02], p = 0.001) compared to the AI baseline. Differences for personality-based predictions were not significant compared to the AI prediction baseline (-0.00, 95% CI [-0.01, 0.00], p = 0.274). No significant interactions were observed between the subscales, prophecy source, and paranormal score.
@@
2.5 Positive AI attitudes increases belief in AI predictions, especially perceived reliability.
@@
Individuals with more positive attitudes towards AI found predictions based on AI more believable. One point increase in AIAS score [41] led to an increase of perceived validity of AI predictions by 0.04 points (95% CI [0.01, 0.06], p = 0.001) on a 7-point scale. The AIAS score ranged from 4 to 40 (Mean = 25.79, SD = 8.68).
@@
The main effect was augmented by an additional 0.03 points for perceived reliability (95% CI [0.02, 0.05], p < 0.001), while the interactions were not significant for perceived personalization (-0.00, 95% CI [-0.01, 0.01], p = 0.941) and usefulness (0.01, 95% CI [-0.00, 0.03], p = 0.071). This suggests that positive attitudes toward AI/trust in AI is more closely related to perceived reliability among all subscales, which is aligned with the findings of studies that explore the relationship between trust/attitude and reliance in AI [42, 43, 44, 45, 46].
@@
The interaction between AI attitude (AIAS) and prediction source showed that the positive impact of positive AI attitudes on perceived validity was mostly reversed for astrology-based predictions (-0.24, 95% CI [-0.41, -0.08], p = 0.004), while the effect did not differ significantly for personality-based predictions (0.04, 95% CI [-0.11, 0.19], p = 0.590). The interaction effects between subscales, prophecy source, and AI attitude score were not statistically significant.
@@
The positive association was observed across other subscales. Compared to perceived validity, perceived personalization increased slightly less, reduced by 0.06 points (95% CI [-0.11, -0.01], p = 0.018), while interaction was not statistically significant for perceived reliabilty (-0.07, 95% CI [-0.14, 0.01], p = 0.068) and perceived usefulness (0.02, 95% CI [-0.04, 0.09], p = 0.475), suggesting that the effects were comparable to perceived validity. The results suggest that when all else is held constant, the more people are interested in the topic, the more likely they will perceive fictitious predictions to be valid, reliable, useful, and personalized.

### v3

# Super-intelligence or Superstition? Exploring Psychological Factors Influencing Belief in AI Predictions about Personal Behavior

@@ Abstract
Could belief in AI predictions be just another form of superstition? This study investigates psychological factors that influence belief in AI predictions about personal behavior, comparing it to belief in astrology- and personality-based predictions. Through an experiment with 238 participants, we examined how cognitive style, paranormal beliefs, AI attitudes, personality traits, and other factors affect perceived validity, reliability, usefulness, and personalization of predictions from different sources. Our findings reveal that belief in AI predictions is positively correlated with belief in predictions based on astrology and personality psychology. Notably, paranormal beliefs and positive attitudes about AI significantly increased perceived validity, reliability, usefulness, and personalization of AI predictions. Conscientiousness was negatively correlated with belief in predictions across all sources, and interest in the prediction topic increased believability across predictions. Surprisingly, we found no evidence that cognitive style has an impact on belief in fictitious AI-generated predictions. These results highlight the "rational superstition" phenomenon in AI, where belief is driven more by mental heuristics and intuition than critical evaluation. This research advances our understanding of the psychology of human-AI interaction, offering insights into designing and promoting AI systems that foster appropriate trust and skepticism, critical for responsible integration in an increasingly AI-driven world.
Keywords Human-AI Interaction  $\cdot$ Psychology  $\cdot$ Cognitive Biases  $\cdot$ Mental Models  $\cdot$ Belief
@@ 1 Introduction
Artificial Intelligence (AI) is often associated with scientific advancement and frequently viewed in opposition to superstition. Yet, despite its foundations in mathematics and computational logic, the public’s engagement with AI may not be as rational or evidence-based as one might expect. In his philosophical paper, Alexander Wilson explores the concept of "rational superstition" as it relates to techno-optimism, challenging traditional ideas of reason and causality. He argues that technological optimism often involves a quasi-superstitious belief in the power of technology to shape the future, which reflects how irrationality can influence our relationship with technological systems [1]. This behavior, where individuals engage with AI predictions in ways that echo quasi-religious beliefs or divination, has been further observed to be paralleling society’s relationship with astrology or religion [2, 3, 4, 5, 6], starkly contrasting with the ideal view of AI as a neutral, rational tool designed to aid human decision-making [7, 8].
Humans rely on a spectrum of cognitive processes, ranging from rapid, intuitive responses (System 1) to slower, reflective reasoning (System 2), both of which can contribute to cognitive biases and mental shortcuts and skew decision-making [9, 10]. The stakes are high when it comes to AI systems—unchecked, this tendency could drive inappropriate reliance on AI in contexts where its limitations and biases could lead to harmful outcomes [11, 12, 13, 14]. The unscrutinized narratives around AI, often evangelized by Silicon Valley executives and popular media [15], further complicate the picture. From claims of superintelligence to exaggerated portrayals about AI either saving or destroying humanity, these narratives can distort public understanding of what AI can and cannot do [16, 17, 18], fostering irrational trust or fear of AI akin to religious or pseudoscientific beliefs [2, 6]. Such polarized views hinder the development of an informed, rational relationship with AI, ultimately posing barriers to its effective and ethical integration into society [1].
Addressing these challenges requires a collective effort from researchers, developers, policymakers, business leaders, and the public. As AI becomes increasingly integrated into everyday life, from recommender systems to personal assistants, fostering critical and informed engagement with AI technologies is essential. Moving beyond simplistic, technology-focused narratives demands a deeper understanding of the psychological and contextual factors shaping people’s perceptions and decisions about AI [19, 20].
This study aims to shed light on the psychological and contextual factors influencing trust and engagement with predictive systems by examining how individuals evaluate predictions about personal behavior from AI, astrology, and personality psychology in the context of financial decision-making. The context is particularly timely given the growing popularity of astrology among Millennials and Gen Z [21, 22, 23], despite its lack of empirical validity [24, 25]. Surveys reveal that a significant portion of Millennials and Gen Z in the U.S. have made financial decisions based on horoscopes [26], fueling the rise of "financial astrology" as a pseudoscientific trend [27, 28].
To explore these dynamics, an experiment was conducted with 238 participants who were presented with fictitious predictions about their personal investing behavior from three sources—AI for its empirical rigor, astrology for its pseudoscientific appeal, and personality psychology as a mix of perceived relatability and scientific grounding. Participants completed fictitious assessments on astrology and personality traits before engaging in a simulated investment game designed to mimic real-world financial decision-making. They then evaluated the perceived validity, reliability, usefulness, and personalization of the predictions. Participants were randomly assigned to receive either positive (N = 119) or negative (N = 119) predictions about their future investment outcomes, allowing us to examine how the valence of predictions influenced their responses.
@@
The study extends prior research by exploring how cognitive style, paranormal beliefs, gullibility, attitudes toward AI, and personality traits influence belief in predictions from these sources. Drawing on insights from work on the effects of user characteristics on the perception and trust of AI-generated advice [29, 30], we consider the interplay between cognitive style and paranormal beliefs, including astrology [31, 32, 33, 34, 35], building on findings about how cognitive style affects belief in (mis)information and conspiracy theories on social media [36, 37] and fake news [38]. Demographic factors such as age, gender, education, and prior experience with AI, astrology, and personality psychology, were also analyzed, alongside participants’ interest in investing behavior.
Guided by this framework, the study tested the following hypotheses:
@@
  • H2: Individuals with a more analytic cognitive style will find AI predictions more credible and reliable than astrology- or personality-based predictions.
  • H3: Demographic factors and personal attributes (cognitive reflection, trust in AI, paranormal beliefs, gullibility, personality, prior experience, level of interest in investing behavior) will moderate the perceived validity, reliability, usefulness, and personalization of predictions from different sources.
Overall, this study examines the psychological and contextual factors that shape trust and engagement with AI systems, offering insights into the "rational superstition" phenomenon and its implications for decision-making in an increasingly AI-driven world. By addressing the interplay between cognitive biases, mental models, and societal narratives, it highlights the need for user-centered design approaches that foster transparency, critical thinking, and accurate mental models. The findings provide actionable strategies for ethical and responsible AI design, public policy, and educational initiatives aimed at promoting AI literacy and informed decision-making. Finally, it underscores the need for interdisciplinary collaboration to design and promote AI systems that align with user needs and societal values, ensuring that AI technologies are integrated responsibly and equitably into everyday life.
@@ 2 Results
Our results with 238 participants demonstrate that belief in AI predictions is positively correlated with belief in astrology, suggesting that non-rational factors, such as superstitious thinking, significantly influence how individuals engage with AI systems. Participants with higher paranormal beliefs and positive attitudes toward AI were more likely to perceive AI predictions as valid, reliable, useful, and personalized, even when faced with fictitious outputs. Surprisingly, cognitive style did not significantly influence belief in AI, astrology, or personality-based predictions. Additionally, conscientiousness was negatively correlated with belief in predictions across all sources, while interest in investing behavior (prediction topic) increased believability. These findings emphasize the importance of developing better mental models and public education to foster critical and rational AI engagement, rather than reliance on speculative thinking.
@@
To challenge the assumption that people approach AI predictions with more rationality than other speculative forms of prediction, we hypothesized that there would be a correlation between people’s perceptions of AI predictions and perceptions of predictions based on astrology and personality (H1). If this were not true, we should see less correlation between the two—people who find AI predictions more believable would find astrology predictions less believable, and vice versa.
To test this hypothesis, a multiple linear regression model was applied to the data in wide format, with believability score for AI predictions as the outcome variable and believability scores for astrology- and personality-based predictions as predictor variables, along with control variables (cognitive style, paranormal beliefs, gullibility, AI attitude/trust in AI, big five personality, familiarity with the prediction sources (AI, astrology, personality), interest in topic of prediction, age, gender, education level). Believability scores for each prediction were calculated by averaging across the four subscales (perceived validity, reliability, usefulness, and personalization), which were drawn from existing literature on the different qualities of a statement that contribute to persuasion, technology adoption, and user attitudes (see Section 7.3.1.
The results showed a significant positive correlation between belief in AI predictions and belief in astrology- and personality-based predictions (see Figure 1), suggesting that people do not engage with AI predictions as rationally as popular narratives might suggest. The multiple linear regression analysis explained a significant proportion of the variance in the AI overall score ($R^{2}=0.7606$, Adjusted $R^{2}=0.7337$, $F(24,213)=28.2$, $p<0.001$). Zodiac overall score (Estimate = 0.3119, $p<0.001$) and the personality overall score (Estimate = 0.4585, $p<0.001$) were significant predictors of the AI overall score, supporting the hypothesis that belief in astrology and personality-based predictions is positively associated with belief in AI predictions.
Figure 1: Relationship between belief in AI predictions and astrology predictions (left) and belief in AI predictions and personality predictions (right), measured on a 7-point Likert scale (1 = Strongly disagree, 7 = Strongly agree). Scatterplots display individual scores with linear regression lines indicating positive correlations, and the shaded area represents the 95% confidence interval for the regression lines.
@@
To examine the influence of cognitive and psychological factors including cognitive style (H2), paranormal beliefs, trust in AI, personality traits, demographic factors, etc. (H3) on belief in AI predictions, a mixed-effects model was fitted to the data in a nested long format, with prediction source (AI, astrology, personality) and believability subscale (perceived validity, reliability, usefulness, and personalization) as categorical predictors, and subscale score as the outcome variable. This approach allowed examination of the main effects and moderating effects of psychological and cognitive factors on belief in AI predictions.
@@
The results indicate that participants generally found fictitious predictions from AI, astrology, and personality sources to be convincing. The baseline belief score, associated with the AI prophecy source, "Validity" subscale, "Positive" prophecy group, "Female" gender, and "Bachelor’s" education level, was 5.10 (p < 0.001) on a 7-point Likert scale. The main effects of prophecy source on the perceived believability of AI predictions revealed that astrology-based predictions were, on average, rated 0.66 points lower in validity compared to AI predictions (95% CI [-0.93, -0.39], p < 0.001). In contrast, the difference for personality-based predictions was not statistically significant (Estimate = 0.07, 95% CI [-0.18, 0.32], p = 0.593), suggesting they were perceived similarly to AI predictions.
@@
There were significant main effects of subscales on the perception of AI predictions. On average, perceived personalization was rated 0.23 point higher than perceived validity (95% CI [0.08, 0.38], p = 0.002). Conversely, perceived reliability was rated 0.91 points lower than perceived validity (95% CI [-1.13, -0.68], p < 0.001), and perceived usefulness was rated 0.63 points lower (95% CI [-0.83, -0.44], p < 0.001). These relationships between subscales and prediction sources are visualized in Figure 2.
The interaction between subscales and prophecy source showed that these main effects were generally consistent across AI, astrology, and personality-based predictions, though some effects were more pronounced for astrology. Specifically, perceived personalization was 0.23 points higher (95% CI [0.02, 0.44], p = 0.035) and perceived reliability was 0.31 points lower (95% CI [-0.62, -0.00], p = 0.048) for astrology-based predictions compared to AI predictions. Other interaction terms were not significant, suggesting that the subscale effects were not substantially different across the three prediction sources.
@@
Additionally, compared to the baseline "Positive" prediction group, the "Negative" prediction group saw a significant decrease in perceived validity by 1.19 points (95% CI [-1.53, -0.85], p < 0.001). No significant interactions between prophecy group and source were found, indicating that this effect was consistent across AI, astrology, and personality predictions. This suggests that while participants generally found fictitious predictions believable, those who received negative predictions were less likely to find them valid.
@@
Figure 2: Boxplot illustrating the distribution of subscale scores (7-point Likert scale: 1 = Strongly disagree, 7 = Strongly agree) across prophecy sources and subscales. The boxes represent the interquartile range, with green triangles indicating the means and notches marking the 95% confidence intervals for the medians.
@@
2.3 There is no evidence of correlation between cognitive style and belief in predictions.
We hypothesized that cognitive style, which has been shown to influence how people process and believe information and make decisions [31, 32, 33, 34, 35, 38], would significantly impact how individuals engage with AI predictions compared to astrology- or personality-based predictions (H2). Specifically, we anticipated a positive association between cognitive style and belief in AI predictions, with negative interactions between cognitive style and prediction sources (astrology and personality) relative to AI, which served as the reference category.
@@
However, the results did not support this hypothesis. The composite cognitive score—derived from the Cognitive Reflection Test (CRT-2) [39] and the Need for Cognition (NFC-6) [40]—did not significantly increase the perceived validity of AI predictions. While a one-point increase in cognitive score was associated with a 0.13-point increase in perceived validity, this effect was not statistically significant (95% CI [-0.01, 0.26], p = 0.065). The cognitive score ranged from -4.29 to 2.61 (Mean = 0.0, SD = 1.44).
Interaction terms further supported the lack of evidence of the association between cognitive style and belief in predictions. Higher cognitive scores were significantly associated with decreased perceived reliability (Estimate = -0.11, 95% CI [-0.20, -0.02], p = 0.021) and usefulness (Estimate = -0.12, 95% CI [-0.20, -0.04], p = 0.004), compared to the reference subscale of validity. No significant association was found with personalization (Estimate = -0.05, 95% CI [-0.11, 0.01], p = 0.092). Interactions between cognitive style and prediction source (relative to AI, the reference category) were negative for astrology (Estimate = -0.10, 95% CI [-0.21, 0.01], p = 0.079) and personality (Estimate = -0.07, 95% CI [-0.18, 0.03], p = 0.170), though neither interaction was statistically significant. Similarly, three-way interactions between subscales, prediction source, and cognitive score were non-significant.
To illustrate these findings, contrast plots in Figure 3 present the fixed effects of the model, including main effects and interactions. The top panel visualizes the relationship between cognitive style and subscale scores (validity, personalization, reliability, and usefulness), highlighting overall trends across evaluation dimensions. The bottom panel shows interactions between cognitive style and prediction sources—AI, astrology, and personality—on subscale scores. Contrast plots are presented for predictors central to the study’s hypotheses to visually highlight their effects. Results for other predictors, which also contribute to the findings but are less central to the main hypotheses, are summarized in writing.
Figure 3: Contrast plots showing the relationship between cognitive style (Range: -4.29 to 2.61; higher values indicate a more analytic style) and centered subscale scores (7-point Likert scale). The top panel illustrates interactions across subscales (validity, personalization, reliability, usefulness), with "Validity" serving as the reference level. The bottom panel shows interactions across prediction sources (AI, astrology, personality), with "AI" as the reference level. Blue lines represent predicted contrasts with 95% confidence intervals, and gray points show individual observations. (*p < 0.05, **p < 0.01, ***p < 0.001). Contrast plots are presented for predictors central to the study’s hypotheses. Results for other predictors are summarized in writing within the Results section.
@@
2.4 Higher paranormal beliefs increase perceived validity, reliability, usefulness, and personalization of AI predictions.
@@
One of the most notable findings was the positive association between paranormal beliefs and belief in AI predictions. Results from the mixed effects model revealed that higher scores on the shortened Revised Paranormal Belief Scale (R-PBS) [41], significantly increased the perceived validity of AI predictions. Specifically, each one-point increase in paranormal beliefs led to an average 0.02-point increase in perceived validity (95% CI [0.01, 0.03], p = 0.001) on a 7-point scale. The paranormal belief scores, centered for analysis, ranged from 15 to 95 (Mean = 45.6, SD = 20.08).
@@
Interactions between paranormal beliefs and subscales showed that this effect was even stronger for perceived reliability (Estimate = 0.01, 95% CI [0.00, 0.01], p = 0.020) and perceived usefulness (Estimate = 0.01, 95% CI [0.01, 0.02], p < 0.001). However, there was no significant effect on perceived personalization (Estimate = 0.00, 95% CI [-0.00, 0.01], p = 0.427), indicating that the impact of paranormal beliefs on perceived personalization did not differ significantly from perceived validity.
@@
Interestingly, paranormal beliefs were an even stronger predictor of belief in astrology-based predictions, which aligns with the nature of the scale, as it includes questions about astrology (see Section 7.3). Each point increase in paranormal beliefs led to an additional 0.01-point increase in perceived validity for astrology predictions (95% CI [0.01, 0.02], p = 0.001) compared to AI. In contrast, differences in personality-based predictions were not significant when compared to AI predictions (Estimate = -0.00, 95% CI [-0.01, 0.00], p = 0.274). No significant interactions were observed between the subscales, prophecy source, and paranormal score. The contrast plots in Figure 4 visualize these results.
Figure 4: Contrast plots showing the relationship between paranormal beliefs (higher values indicate stronger paranormal beliefs) and centered subscale scores (7-point Likert scale). Top: interactions across subscales (validity [reference level], personalization, reliability, usefulness). Bottom: interactions across prediction sources (AI [reference level], astrology, personality). Blue lines represent predicted contrasts with 95% confidence intervals, and gray points show individual observations. The plot ratio was adjusted to align with other contrast plots for visual comparability. (*p < 0.05, **p < 0.01, ***p < 0.001)
@@
2.5 Positive attitudes toward AI increase belief in AI predictions, especially perceived reliability.
@@
Individuals with more positive attitudes towards AI found AI-based predictions more believable. Each one-point increase in the AI Attitude Scale (AIAS) [42] was associated with a 0.04-point increase in the perceived validity of AI predictions (95% CI [0.01, 0.06], p = 0.001) on a 7-point scale. The AIAS score ranged from 4 to 40 (Mean = 25.79, SD = 8.68).
@@
This effect was particularly pronounced for perceived reliability, which increased by an additional 0.03 points (95% CI [0.02, 0.05], p < 0.001), while the interactions were not significant for perceived usefulness (Estimate = 0.01, 95% CI [-0.00, 0.03], p = 0.071) or perceived personalization (Estimate = -0.00, 95% CI [-0.01, 0.01], p = 0.941). This suggests that positive attitudes towards AI are most strongly associated with higher perceived reliability, aligning with previous studies on the relationship between trust in AI and reliance on its predictions [43, 44, 45, 46, 47].
@@
The interaction between AI attitudes and prediction source revealed that the positive impact of positive AI attitudes on perceived validity was largely reversed for astrology-based predictions, which saw a 0.24-point decrease (95% CI [-0.41, -0.08], p = 0.004). For personality-based predictions, the effect did not differ significantly from AI predictions (Estimate = 0.04, 95% CI [-0.11, 0.19], p = 0.590). There were no statistically significant interaction effects between subscales, prophecy source, and AI attitude score. These findings are visualized in the contrast plots in Figure 5.
Figure 5: Contrast plots showing the relationship between attitude toward AI (higher values indicate more positive attitudes) and centered subscale scores (7-point Likert scale). Top: interactions across subscales (validity [reference level], personalization, reliability, usefulness). Bottom: interactions across prediction sources (AI [reference level], astrology, personality). Blue lines represent predicted contrasts with 95% confidence intervals, and gray points show individual observations. The plot ratio was adjusted to align with other contrast plots for visual comparability. (*p < 0.05, **p < 0.01, ***p < 0.001)
@@
The positive association was observed across other subscales. Compared to perceived validity, perceived personalization increased slightly less, reduced by 0.06 points (95% CI [-0.11, -0.01], p = 0.018), while interaction was not statistically significant for perceived reliabilty (Estimate = -0.07, 95% CI [-0.14, 0.01], p = 0.068) and perceived usefulness (Estimate = 0.02, 95% CI [-0.04, 0.09], p = 0.475), suggesting that the effects were comparable to perceived validity. The results suggest that when all else is held constant, the more people are interested in the topic, the more likely they will perceive fictitious predictions to be valid, reliable, useful, and personalized.